OpenStudy (sushi121212):

Find the zeros of the function: f(x) x^2-4x√5 +19

OpenStudy (sushi121212):

I don't understand how to solve it with the 4x√5. Do I have to get rid of it?

OpenStudy (jango_in_dtown):

x^2-4x√5 +19=0 implies x^2-2x.2√5+(2√5)^2+19-(2√5)^2=0 implies (x-2√5)^2=1 or, x-2√5=1,-1 or, x=1+2√5,-1+2√5

OpenStudy (sushi121212):

Okay, thank you!!

OpenStudy (sushi121212):

Actually, I can't really tell what you're writing. Can you write it again? I'm sorry...